Marviw is 5 years older then her brother. 5 years ago she was 2 times older then her brother. how old is she now?
Answer by ewatrrr(24785)


Hi
Let x and (x+5) represent the brother's and Maria ages NOW respectively
Note: Maria 'is, was, and will always be' 5 years older than her brother.
5 years ago...***
(x+5) - 5 = 2(x-5)
Solving for x
x = 2x - 10
10 = x , bother's age NOW. Maria is 15
CHECKING our Answer***
10 = 2*5